Required information

Skip to question


[The following information applies to the questions displayed below.]

 

Causwell Company began 2021 with 12,000 units of inventory on hand. The cost of each unit was $5.00. During 2021 an additional 32,000 units were purchased at a single unit cost, and 22,000 units remained on hand at the end of 2021 (22,000 units therefore were sold during 2021). Causwell uses a periodic inventory system. Cost of goods sold for 2021, applying the average cost method, is $118,800. The company is interested in determining what cost of goods sold would have been if the FIFO or LIFO methods were used.


Required:
1. Determine the cost of goods sold for 2021 using the FIFO method. [Hint: Determine the cost per unit of 2021 purchases.] (Do not round intermediate calculations.)


Required:
2. Determine the cost of goods sold for 2021 using the LIFO method. (Do not round intermediate calculations.)
Answer #1

req 1) 

step 1) 12,000+ 32,000= 44,000 units available for sale

step 2) 12,000+32,000-22,000= 22,000

step 3) COGS/ units sold:  118,000/22,000= 5.44

step 4) 44,000*5.4= 237,600

step 5) 12,000* 5units= 60,000

step 6) 237,600-60,000= 177,600

step 7) 177,600/32,000= 5.55

step 8) 22,000-12,000= 10,000

step 9) 10,000*5.55= 55,500

step 10) 60,000+55,500= 115,500 Answer


Req 2)

22,000* 5.55= 122,100 answer

answered by: anonymous